1. The third and the six term of a G.P are 108 and -32 respectively find the sum of first seven term.
2. The first and last term of a G.P are 2048 respectively the sum of the series is 2730. Find the number of terms and common ratio.
1
Expert's answer
2020-02-14T10:31:12-0500

1. Write the third and sixth terms of geometric progression:


"b_3=b_1q^{3-1}=108,\\
b_6=b_1q^{6-1}=-32.\\"

Solve that system. Divide one by another:


"\frac{108}{-32}=\frac{q^2}{q^5},\\
\space\\
q^3=-\frac{8}{27},\\
\space\\
q=-\sqrt[3]\frac{8}{27}=-\frac{2}{3}."

"b_1=\frac{108}{q^2}=243."

Sum of the first seven terms:


"S_7=\frac{b_1(q^7-1)}{q-1}=\frac{243[(-2/3)^7-1]}{-2/3-1}=\frac{463}{3}."

2. This is a "reverse" problem for a sequence with "b_1=b_n=2048" and "S_n=2730":

"S_n=\frac{2048(q^n-1)}{q-1}=2730."

Since the first and the last terms are equal, the common ratio (that makes them equal) is 1. Since such a value of a common ratio results in 0 in the denominator of the sum equation, the problem can't be solved.